how do you solve this equations m4+7m3+12m2=0
Answer by checkley77(12844)

m^4+7m^3+12m^2=0 FIRST WE EXTRACT THE M^2 TERM.
M^2(M^2+7M+12)=0 NOW WE FACTOR THE QUADRATIC EQUATION.
M^2(M+4)(M+3)=0 NOW WE SET EACH TERM EQUAL TO ZERO.
M^2=O
M=0 ANS.
M+4=0
M=-4 ANS.
M+3=0
M=-3 ANS.